# Inverse k-visibility for RSSI-based Indoor Geometric Mapping

### v1

@@ VI-D Experimental Discussion Between Trajectories
This paper presents an innovative approach to Wifi-based geometric mapping, which effectively estimates the approximate layout of indoor spaces, especially the free space. These free spaces are necessary for the robot to plan paths and explore unfamiliar environments. Notably, our method does not depend on crowdsourcing or exteroceptive sensors such as a camera, radar, or lidar. Further, Our results demonstrate an overall increase in mapping accuracy by approximately 6.34% on average with about 85% fewer data points required, compared to using a single router setup.
@@
As for future work, it will be necessary to investigate machine learning techniques to enhance the quality of the maps during post-processing. Combining WiFi-based localization for precise trajectory tracking with our WiFi-based mapping will provide a more comprehensive autonomous WiFiSLAM system. Additionally, strategies for active trajectory planning to improve map accuracy will be explored. Further research should also focus on developing advanced filtering systems that account for distance in the case of the large-scale map. Due to the characteristics of the RSSI signal, multiple routers should be incorporated for a large-scale environment. Strategic router placement is also needed for optimizing signal strength distribution and improving data quality. Routers should be positioned with minimal signal overlap and reduce areas of weak or no coverage. In most cases, this can be achieved by placing the routers at key locations such as the corners and central points of large rooms, while ensuring line-of-sight from any points.
@@ Acknowledgements
We would like to thank Jill Aghyourli Zalat for her early work formulating the SfW concepts and Ishaan Mehta for his feedback and assistance with data acquisition. This research is supported by Natural Sciences and Engineering Research Council of Canada (NSERC).

### v2

@@ VIII Future work and Conclusions
This paper presents an innovative approach to WiFi-based geometric mapping, which effectively estimates the approximate layout of indoor spaces, especially the free space. These free spaces are necessary for the robot to plan paths and explore unfamiliar environments. Notably, our method does not depend on crowdsourcing or exteroceptive sensors such as a camera, radar, or LiDAR. Further, Our results demonstrate an overall increase in mapping accuracy by approximately 6.34% on average, with about 85% fewer data points required, compared to using a single router setup.
As for future work, it will be necessary to investigate machine learning techniques to enhance the quality of the maps during post-processing. Combining WiFi-based localization for precise trajectory tracking with our WiFi-based mapping will provide a more comprehensive autonomous WiFiSLAM system. Additionally, strategies for active trajectory planning to improve map accuracy will be explored, including frontier-based exploration, which is navigating to frontiers between known free and unknown space, and information-gain policies, which is choosing next poses that maximize expected reduction in map uncertainty [yamauchi1997frontier, bourgault2002information, stachniss2005information, charrow2015information, bircher2016receding].
More recent active SLAM methods go further by explicitly planning actions to reduce uncertainty in the SLAM belief. Approaches based on belief-space planning evaluate candidate trajectories by their expected reduction in pose/map uncertainty rather than just area coverage [vallve2015active, ahmed2023active]. In parallel, loop-closure aware planning explicitly drives the robot to re-observe past areas from new perspective to correct drift and refine the map [stachniss2004exploration]. By deliberately revisiting and closing loops during exploration, the robot can substantially lower its localization error and produce a more consistent map than purely frontier-driven strategies.
To eliminate residual “unknown” bands that can arise between router dominance zones (e.g., Fig. 10-(d)), we recommend inserting short cross-connecting passes. These passes are implemented as brief perpendicular trajectories intersecting the main routes, directly traversing ambiguous or previously unseen strips between WiFi router dominance zones. By providing additional kk=0 classification (line-of-sight) observations in these transitional areas, such passes resolve uncertain boundaries and eliminate “unknown” bands[choset2001coverage, galceran2013survey]. This targeted strategy improves overall map completeness by ensuring that previously unmapped free-space segments are classified and integrated into the map.
@@
Further research should also focus on developing advanced (multipath-aware filtering systems that account for distance in the case of the large-scale map. Due to the characteristics of the RSSI signal, multiple routers should be incorporated for a large-scale environment, preferably across multiple bands (e.g., 2.4/5 GHz) to increase robustness. Strategic router placement is also needed for optimizing signal strength distribution and improving data quality. Routers should be positioned with minimal signal overlap and reduce areas of weak or no coverage. In most cases, this can be achieved by placing the routers at key locations such as the corners and central points of large rooms, while ensuring line-of-sight from any point.
@@ Acknowledgements
We would like to thank Jill Aghyourli Zalat for her early work formulating the SfW concepts, and Ishaan Mehta for his feedback and assistance with data acquisition. This research is supported by the Natural Sciences and Engineering Research Council of Canada (NSERC).